OpenStudy (anonymous):

how do you put this cos^2(pi/11)-sin^2(pi/11) in double-angle formula?

OpenStudy (campbell_st):

well its cos(pi/11 + pi/11)

OpenStudy (campbell_st):

because cos(2A) = cos(A +A) = cos(A) cos(A) - Sin^(A)sin^(A) = cos^2(A) - sin^(A)

OpenStudy (anonymous):

Thank you! What is an angle between 0 and 2pi that is coterminal with -5pi/6?